# sample.bib

### v2

@@ Abstract
Medical image segmentation takes an important position in various clinical applications. 2.5D-based segmentation models bridge the computational efficiency of 2D-based models with the spatial perception capabilities of 3D-based models. However, existing 2.5D-based models primarily adopt a single encoder to extract features of target and neighborhood slices, failing to effectively fuse inter-slice information, resulting in suboptimal segmentation performance. In this study, a novel momentum encoder-based inter-slice fusion transformer (MOSformer) is proposed to overcome this issue by leveraging inter-slice information at multi-scale feature maps extracted by different encoders. Specifically, dual encoders are employed to enhance feature distinguishability among different slices. One of the encoders is moving-averaged to maintain consistent slice representations. Moreover, an inter-slice fusion transformer (IF-Trans) module is developed to fuse inter-slice multi-scale features. The MOSformer is evaluated on three benchmark datasets (Synapse, ACDC, and AMOS), achieving a new state-of-the-art with 85.63%, 92.19%, and 85.43% DSC, respectively. These results demonstrate MOSformer’s competitiveness in medical image segmentation.
@@ 1 Introduction
Medical image segmentation plays a crucial role in numerous clinical applications, such as computer-aided diagnoses [antonelli2022medical], image-guided interventions [han2024respiratory][huang2025real][chen2025gvm], and surgical robotics [zhang2024ai][li2023three][zhou2022learning]. UNet [ronneberger2015u] and its variants [zhou2019unet++][huang2020unet][huang2024spironet] have been widely used in this field, achieving tremendous success in different medical imaging modalities. However, accurate and efficient segmentation of 3D medical images still remains a non-trivial task [tajbakhsh2020embracing].
Current mainstream segmentation methods can be classified into two categories: 2D-based and 3D-based methods [zhang2022bridging]. 2D-based methods split 3D images into 2D slices and segment them individually, while 3D-based methods divide 3D images into smaller patches and then segment these patches individually. Despite impressive performance achieved by state-of-the-art methods [azad2024medical], they still exhibit some limitations. Most 2D-based methods focus on architecture design to enhance intra-slice representations for better performance, such as incorporating attention modules [mou2021cs2][roy2022lwmla] or adopting transformers [cao2022swin][chen2024transunet]. However, these methods overlook inter-slice cues, which is also crucial for accurate segmentation. In contrast, 3D-based methods can capture intra- and inter-slice information for segmentation but demand substantial GPU memory and computational resources. Additionally, they tend to perform poorly in images with anisotropic voxel spacing since they are primarily designed for 3D images with nearly isotropic voxel spacing [cciccek20163d][yu2017automatic].
In order to combine advantages of 2D-based and 3D-based methods, some studies have been done to explore 2.5D-based segmentation models [zhang2022bridging]. The main idea of these methods is to fuse inter-slice (neighborhood slices) information into 2D-based models when segment specific slices (target slices). The most direct way to achieve inter-slice fusion is by concatenating slices as multi-channel inputs. However, it is inefficient, making models challenging to extract useful features for target slices [zhang2022bridging]. Therefore, some studies focus on exploring “smart” ways of inter-slice fusion. Most of them formulate 2D slices as time sequences and adopt recurrent neural network (RNN) [chen2016combine], transformers [yan2022after][hung2022cat] or attention mechanisms [zhang2020sau] to fuse inter-slice information.
While current 2.5D-based methods have achieved impressive segmentation results, they struggle with distinguishing individual slices during inter-slice fusion, and consequently fail to learn reliable inter-slice representations essential for accurate segmentation [zhang2022bridging]. The root of this problem lies in the use of a single encoder to process all input slices, resulting in the same feature distributions across the feature space, as shown in Fig. 1 (a). For example, the features of the $i$-th slice remain identical whether it is considered as the target slice or the neighboring slice. This indistinguishability becomes problematic in scenarios where consecutive slices, such as the $i$-th and the $(i+1)$-th, are target slices, respectively. Models fail to differentiate the $i$-th slice’s features as belonging to the target or the neighboring slice, thereby hampering the extraction of valuable inter-slice information for segmentation.
To address the above issue, a novel 2.5D-based segmentation model, MOSformer, MOmentum encoder-based inter-Slice fusion transformer is proposed to effectively leverage inter-slice information for 3D medical image segmentation. MOSformer follows the design of the U-shape architecture [ronneberger2015u]. In order to enhance feature distinguishability of each slice, dual encoders are utilized in our model. One for target slices and the other for neighborhood slices. Parameters of the target slice encoder are updated by back-propagation, and parameters of the neighborhood slice encoder are updated using a momentum update. Therefore, features can remain distinguishable and consistent, promoting inter-slice fusion, as shown in Fig. 1 (b). Furthermore, an efficient inter-slice fusion transformer (IF-Trans) module is proposed to capture inter-slice cues at multi-scale feature maps based on Swin transformer [liu2021swin].
@@ 2 Related Works
Several 2.5D-based approaches have been proposed for efficient medical image segmentation by leveraging inter-slice information. Early methods concatenated multiple consecutive 2D slices into a multi-channel input and adopted 2D-based models to segment specific regions of the middle slice [yu2018recurrent][li2021acenet]. However, 2D multi-channel inputs make models difficult to extract information of individual information [zhang2022bridging], limiting the potential of 2.5D-based model. To solve the above problem, some works treated continuous 2D slices as temporal sequences and utilized recurrent neural networks (RNNs) [chen2016combine][yang2018towards] to learn inter-slice information. For example, Chen et al. [chen2016combine] suggested a 2.5D segmentation framework combining with $k$-UNet and bi-directional convolutional LSTM (BDC-LSTM) to integrate inter-slice information. Although RNNs can help improve the performance of 2.5D-based models to some extent, training costs of these methods are considerably high [pang2019deep]. Instead of RNNs, recent studies utilized attention mechanisms or transformers to fuse inter-slice information at feature level effectively. Zhang et al. [zhang2020sau] proposed an attention fusion module to refine segmentation results by fusing the information of adjacent slices. Li et al. [li2021learning] adopted a 2.5D coarse-to-fine architecture, which benefits from the inter-slice context knowledge from consistency context similarity and discrepancy context. Guo et al. [guo2021transformer] adopted 2D UNet as the backbone and fused inter-slice information via a transformer at the bottom layer of the encoder. Yan et al. [yan2022after] proposed AFTer-UNet with an axial fusion mechanism based on transformer to fuse intra- and inter-slice contextual information. Hung et al. [hung2022cat][hung2024csam] introduced novel cross-slice attention mechanisms based on transformer to learn cross-slice information at multiple scales. However, the aforementioned methods fail to capture useful inter-slice information for the target slice which needed to be segmented, since they use a single encoder to extract slice features, making models difficult to distinct target slices and neighborhood slices [zhang2022bridging].
Different from previous works, we adopt dual encoders with a momentum update to extract features of target slices and neighborhood slices, respectively. We demonstrate such design can make features of target slices and neighborhood slices distinguishable and consistent, further boosting inter-slice fusion.
@@ 2.2 Transformers in Medical Image Segmentation
Recently, with the tremendous success of vision transformer (ViT) [dosovitskiy2020image] in various computer vision tasks [carion2020end][lijin2024dual], many works have explored using transformers in medical image segmentation. Compared with CNNs, transformers can capture long-range dependencies by sequence modeling and multihead self-attention (MHSA) [dosovitskiy2020image], achieving better segmentation performance. Chen et al. [chen2024transunet] proposed TransUNet, combining UNet [ronneberger2015u] and transformer, where transformer encodes feature maps from CNN encoder to extract global contexts for the decoder to generate segmentation results. Cao et al. [cao2022swin] are the first to employ a pure transformer architecture for medical image segmentation.
Figure 2: The architecture of MOSformer. It consists of dual encoders: the momentum encoder for extracting neighborhood slice features and the encoder for extracting target slice features. The IF-Trans is designed to achieve inter-slice fusion at multi-scale features. After that, fused features are fed to the CNN decoder to generate segmentation maps of target slices.
Convolutional layers in UNet are all replaced by blocks in Swin transformer [liu2021swin]. However, this architecture does not obtain better performance [you2022class]. Huang et al. introduced MISSformer [huang2023missformer], which incorporates an encoder-decoder architecture built on enhanced transformer blocks. These blocks are connected through the ReMixed transformer context bridge, enhancing the model’s ability to capture discriminative details. You et al. [you2022class] presented CASTformer with a class-aware transformer module to better capture discriminative regions of target objects. Moreover, they utilized adversarial learning to boost segmentation accuracies. However, the 2D-based methods mentioned above face limitations in leveraging inter-slice information, which hinders their potential for further performance improvements. Some attempts have been made to build 3D-based transformer segmentation models. UNETR [hatamizadeh2022unetr] pioneered the use of a transformer-based encoder to learn global contexts from volumetric data. CoTr [xie2021cotr] introduced a deformable self-attention mechanism to reduce computational complexity. However, simplifying self-attention may cause contextual information loss [yan2022after]. nnFormer [zhou2023nnformer] is an interleave architecture, where convolution layers encode precise spatial information and transformer layers fully explore global dependencies. Similar to Swin transformer [liu2021swin], a computationally-efficient way to calculate self-attention is proposed in nnFormer.
In this work, we proposed a inter-slice fusion transformer module based on Swin transformer [liu2021swin] to effectively fuse inter-slice information. Unlike current methods only use transformer to learn intra-slice contexts, we extent self-attention of transformers to explore both intra- and inter-slice cues.
@@ 3.1 Overall Architecture
The detailed architecture of MOSformer is shown in Fig. 2. Like most previous works for medical image segmentation, we utilize a hybrid encoder-decoder architecture, combining the advantages of CNNs and transformers [zhou2023nnformer]. $\bm{x}_{i}\in\mathbb{R}^{C\times H\times W}$ is the input of the encoder and represents the target slice for segmentation, where $i$ indicates the $i$-th slice of a 3D volume $\bm{X}\in\mathbb{R}^{C\times H\times W\times D}$, where $C$, $H$, $W$, and $D$ denote the channel, height, width, and depth of $\bm{X}$. $\bm{x}_{j}\in\mathbb{R}^{C\times H\times W}$ are inputs of the momentum encoder and represent neighborhood slices of $\bm{x}_{i}$, $j\in\left[i-s,i+s\right]/\left\{i\right\}$ and $s$ represents the $s$-th neighborhood of $\bm{x}_{i}$. Finally, the model generates the segmentation map $\bm{y}_{i}\in\mathbb{R}^{C_{0}\times H\times W}$ of $\bm{x}_{i}$, where $C_{0}$ is the number of label classes.
We utilize ResNet50 [he2016deep] as our encoder to extract multi-scale features of input slices. Dual encoders with a momentum update are adopted in MOSformer to strengthen feature distinguishbility and maintain feature consistency. Furthermore, IF-Trans modules are used at multi-scale ($1/2$, $1/4$, $1/8$, and $1/16$) to fuse fuse inter-slice features extracted by dual encoders. Then the fused features are sent to the decoder via skip connections. The final segmentation predictions are derived via a segmentation head ($1\times 1$ convolutional layer).
@@ 3.2 Dual Encoders with A Momentum Update
Conventional 2.5D-based methods utilize a single encoder to process all input slices and then fuse inter-slice information at feature level. This may make models chanllenging to distinguish each slice [zhang2022bridging] since they are from the same feature space, as illustrated in Fig. 1 (a). Models cannot focus on capturing discriminative inter-slice representations for target slices. A simple idea is to use two independently updated encoders to process neighborhood slices and target slices, respectively. However, this approach achieves suboptimal performance in experiments conducted in Section 5.2. We hypothesize that such unsatisfactory performance is caused by two independently update encoders that reduce slice features’ consistency which is also crucial for inter-slice fusion.
Inspired by [he2020momentum], a momentum update approach is adopted to overcome the above issue. Parameters of target slice encoder $\bm{\theta}_{1}$ are updated by standard back-propagation. Parameters of neighborhood slice encoder $\bm{\theta}_{2}$ are updated by:
@@
where $m\in\left[0,1\right)$ is a momentum coefficient ($0.1$ in default). We investigate the effect of $m$ in Section 5.3 and conclude that $m$ should be relatively small to make features consistent. Under this circumstance, slice features are distinguishable and consistent, as shown in Fig. 1 (b) and Fig. 7 (c), facilitating the model in effectively capturing inter-slice information.

### v3

# Data curation, Methodology, Writing - original draft

@@ Abstract
Medical image segmentation takes an important position in various clinical applications. 2.5D-based segmentation models bridge the computational efficiency of 2D-based models with the spatial perception capabilities of 3D-based models. However, existing 2.5D-based models primarily adopt a single encoder to extract features of target and neighborhood slices, failing to effectively fuse inter-slice information, resulting in suboptimal segmentation performance. In this study, a novel momentum encoder-based inter-slice fusion transformer (MOSformer) is proposed to overcome this issue by leveraging inter-slice information from multi-scale feature maps extracted by different encoders. Specifically, dual encoders are employed to enhance feature distinguishability among different slices. One of the encoders is moving-averaged to maintain consistent slice representations. Moreover, an inter-slice fusion transformer (IF-Trans) module is developed to fuse inter-slice multi-scale features. MOSformer is evaluated on three benchmark datasets (Synapse, ACDC, and AMOS), achieving a new state-of-the-art with 85.63%, 92.19%, and 85.43% DSC, respectively. These results demonstrate MOSformer’s competitiveness in medical image segmentation.
keywords:
[…]
@@ 1 Introduction
Medical image segmentation plays a crucial role in numerous clinical applications, such as computer-aided diagnoses [1][2], image-guided interventions [3][4], [5][6][7], and surgical robotics [8][9][10][11]. UNet [12] and its variants [13][14][15][16] have been widely used in this field, achieving tremendous success in different medical imaging modalities. However, accurate and efficient segmentation of 3D medical images still remains a non-trivial task [17].
Current mainstream segmentation methods can be classified into two categories: 2D-based and 3D-based methods [18]. 2D-based methods split 3D images into 2D slices and segment them individually, while 3D-based methods divide 3D images into smaller patches and then segment these patches individually. Despite impressive performance achieved by state-of-the-art methods [19], they still exhibit some limitations. Most 2D-based methods focus on architecture design to enhance intra-slice representations for better performance, such as incorporating attention modules [20][21] or adopting transformers [22][23]. However, these methods overlook inter-slice cues, which are also crucial for accurate segmentation. In contrast, 3D-based methods can capture intra- and inter-slice information for segmentation but demand substantial GPU memory and computational resources. Additionally, they tend to perform poorly in images with anisotropic voxel spacing since they are primarily designed for 3D images with nearly isotropic voxel spacing [24][25].
In order to combine the advantages of 2D-based and 3D-based methods, some studies have been done to explore 2.5D-based segmentation models [18]. The main idea of these methods is to fuse inter-slice (neighborhood slices) information into 2D-based models when segmenting specific slices (target slices). The most direct way to achieve inter-slice fusion is by concatenating slices as multi-channel inputs. However, it is inefficient, making it challenging for models to extract useful features for the target slice [18]. Therefore, some studies focus on exploring “smart” ways of inter-slice fusion. Most of them formulate 2D slices as time sequences and adopt recurrent neural network (RNN) [26], transformers [27][28] or attention mechanisms [29] to fuse inter-slice information.
While current 2.5D-based methods have achieved impressive segmentation results, they struggle with distinguishing individual slices during inter-slice fusion, and consequently fail to learn reliable inter-slice representations essential for accurate segmentation [18]. The root of this problem lies in the use of a single encoder to process all input slices, resulting in the same feature distributions across the feature space, as shown in Fig. 1 (a). For example, the features of the $i$-th slice remain identical whether it is considered as the target slice or the neighborhood slice. This indistinguishability becomes problematic in scenarios where consecutive slices, such as the $i$-th and the $(i+1)$-th, are target slices, respectively. Models fail to differentiate the $i$-th slice’s features as belonging to the target or the neighborhood slice, thereby hampering the extraction of valuable inter-slice information for segmentation.
@@
To address the above issue, a novel 2.5D-based segmentation model, MOSformer, MOmentum encoder-based inter-Slice fusion transformer is proposed to effectively leverage inter-slice information for 3D medical image segmentation. MOSformer follows the design of the U-shaped architecture [12]. In order to enhance feature distinguishability of each slice, dual encoders are utilized in our model, with one for target slices and the other for neighborhood slices. Parameters of the target slice encoder are updated by back-propagation, and parameters of the neighborhood slice encoder are updated using a momentum update. Therefore, features can remain distinguishable and consistent, promoting inter-slice fusion, as shown in Fig. 1 (b). Furthermore, we propose an efficient inter-slice fusion Transformer (IF-Trans) that captures inter-slice cues from multi-scale feature maps at each scale, built upon Swin Transformer [30].
@@ 2 Related Works
Figure 2: The architecture of MOSformer. It comprises dual encoders: a momentum encoder that extracts features of the neighborhood slice and an encoder that extracts features from the target slice. IF-Trans is designed to perform inter-slice fusion independently at different scales. The fused features are then fed into a CNN decoder to produce segmentation maps for the target slices. Cylinders in yellow, pink, blue, green, and purple denote feature maps produced by the momentum encoder, the encoder, the IF-Trans, the upsampling operators, and the decoder blocks, i.e.,(Conv+BN+ReLU) * 2, respectively.
@@
Several 2.5D-based approaches have been proposed for efficient medical image segmentation by leveraging inter-slice information. Early methods concatenated multiple consecutive 2D slices into a multi-channel input and adopted 2D-based models to segment specific regions of the middle slice [31][32]. However, concatenating 2D slices as multi-channel inputs hinders the model’s capacity to disentangle and learn slice-specific features [18], thereby constraining the performance of 2.5D-based models. To solve the above problem, some works treated continuous 2D slices as temporal sequences and utilized recurrent neural networks (RNNs) [26][33] to learn inter-slice information. For example, Chen et al. [26] introduced a 2.5D segmentation framework that combines $k$-UNet and bi-directional convolutional LSTM (BDC-LSTM) to integrate inter-slice information. Although RNNs can help improve the performance of 2.5D-based models to some extent, training costs of these methods are considerably high [34]. Instead of RNNs, recent studies utilized attention mechanisms or transformers to fuse inter-slice information at the feature level effectively. Zhang et al. [29] proposed an attention fusion module to refine segmentation results by fusing the information of adjacent slices. Li et al. [35] employed a 2.5D coarse-to-fine architecture that leveraged inter-slice prediction discrepancies as spatial attention cues to refine the initial segmentation. Guo et al. [36] adopted 2D UNet as the backbone and fused inter-slice information via a transformer at the bottom layer of the encoder. Yan et al. [27] proposed AFTer-UNet with an axial fusion mechanism based on transformer to fuse intra- and inter-slice contextual information. Hung et al. [28][37] and Kumar et al. [38] introduced novel cross-slice attention mechanisms based on transformer to learn cross-slice information at multiple scales. However, the aforementioned methods fail to capture useful inter-slice information for the target slice which needs to be segmented, since they use a single encoder to extract slice features, making it difficult for models to distinguish target slices from neighborhood slices [18].
Different from previous works, we adopt dual encoders with a momentum update to extract features of target slices and neighborhood slices, respectively. We demonstrate that such a design can make features of target slices and neighborhood slices distinguishable and consistent, further boosting inter-slice fusion.
@@ 2.2 Transformers in Medical Image Segmentation
Recently, with the tremendous success of vision transformer (ViT) [39] in various computer vision tasks [40][41], many works have explored using transformers in medical image segmentation. Compared with CNNs, transformers can capture long-range dependencies by sequence modeling and multihead self-attention (MHSA) [39], achieving better segmentation performance. Chen et al. [23] proposed TransUNet, combining UNet [12] and transformer, where the transformer encodes feature maps from the CNN encoder to extract global contexts for the decoder to generate segmentation results. Cao et al. [22] are the first to employ a pure transformer architecture for medical image segmentation. Convolutional layers in UNet are all replaced by blocks in Swin transformer [30]. However, this architecture does not achieve better performance [42]. Huang et al. introduced MISSformer [43], which incorporates an encoder-decoder architecture built on enhanced transformer blocks. These blocks are connected through the ReMixed transformer context bridge, enhancing the model’s ability to capture discriminative details. You et al. [42] presented CASTformer with a class-aware transformer module to better capture discriminative regions of target objects. Moreover, they utilized adversarial learning to boost segmentation accuracies. However, the 2D-based methods mentioned above face limitations in leveraging inter-slice information, which hinders their potential for further performance improvements. Some attempts have been made to build 3D-based transformer segmentation models. UNETR [44] pioneered the use of a transformer-based encoder to learn global contexts from volumetric data. CoTr [45] introduced a deformable self-attention mechanism to reduce computational complexity. However, simplifying self-attention may cause contextual information loss [27]. nnFormer [46] is an interleaved architecture, where convolution layers encode precise spatial information and transformer layers fully explore global dependencies. Similar to Swin transformer [30], a computationally efficient way to calculate self-attention is proposed in nnFormer.
In this work, we introduce the inter-slice fusion transformer (IF-Trans), which extends Swin Transformer’s (shifted) window multi-head self-attention [30], (S)W-MSA, to fuse inter-slice information. Unlike prior methods that restrict transformer attention to each 2D slice, IF-Trans lifts (S)W-MSA into the inter-slice domain. Each window attends not only to patches within its own slice but also to the corresponding windows in neighborhood slices. This design preserves Swin Transformer’s computational efficiency while jointly modeling intra-slice and inter-slice context, improving 3D medical image segmentation.
@@ 3.1 Overall Architecture
The detailed architecture of MOSformer is shown in Fig. 2. Like most previous works for medical image segmentation, we utilize a hybrid encoder-decoder architecture, combining the advantages of CNNs and transformers [46]. $\bm{x}_{i}\in\mathbb{R}^{C\times H\times W}$ is the input of the encoder and represents the target slice for segmentation, where $i$ indicates the $i$-th slice of a 3D volume $\bm{X}\in\mathbb{R}^{C\times H\times W\times D}$, where $C$, $H$, $W$, and $D$ denote the channels, height, width, and depth of $\bm{X}$. $\bm{x}_{j}\in\mathbb{R}^{C\times H\times W}$ denotes an input to the momentum encoder corresponding to a neighborhood slice of $\bm{x}_{i}$, where $j\in\left[i-s,i+s\right]\setminus\left\{i\right\}$. The hyperparameter $s$ represents the $s$-th neighborhood of $\bm{x}_{i}$. Finally, the model generates the segmentation map $\bm{y}_{i}\in\mathbb{R}^{C_{0}\times H\times W}$ of $\bm{x}_{i}$, where $C_{0}$ is the number of label classes.
We adopt a lightly modified ResNet-50 [47] to implement the encoders, which extract multi-scale features from input slices. Concretely, the final downsampling stage is replaced with a non-downsampling stage to preserve spatial resolution. Dual encoders with a momentum update are adopted in MOSformer to strengthen feature distinguishability and maintain feature consistency. Furthermore, IF-Trans modules are used at multi-scale ($1/2$, $1/4$, $1/8$, and $1/16$) to fuse inter-slice features extracted by dual encoders (details are provided in Section 3.3). Then the fused features are sent to the decoder via skip connections. The final segmentation predictions are derived via a segmentation head ($1\times 1$ convolutional layer).
@@ 3.2 Dual Encoders with A Momentum Update
Conventional 2.5D-based methods employ a single encoder to process both the target slice and its neighborhood slices, then fuse their features at a later stage. However, because all slices share the same feature space, the model struggles to distinguish target-specific cues from neighborhood context [18], as illustrated in Fig. 1 (a). Consequently, inter-slice fusion may be suboptimal, and fine-grained spatial cues of the target slice can be lost. An intuitive solution is to use two separate encoders to process neighborhood slices and target slices, respectively. In practice, however, updating the parameters of these encoders independently during training leads to inconsistent feature distributions, which again hampers effective fusion.
Figure 3: Schematic of inter-slice fusion transformer (IF-Trans) module. The neighborhood slice number is set to $1$ in this figure, consistent with our default model configuration. It has two successive IF-Trans with different window partitioning configurations. The colored circles indicate feature pixels. The window-based self-attention is expanded to the inter-slice dimension, promoting target slice feature pixels to learn intra- and inter-slice contexts. The black and red arrows denote the data flow within the IF-Trans Module and the fusion process of the IF-Trans Module.
To address this, we draw inspiration from the momentum contrast framework (MoCo) [48] and introduce a momentum encoder for the neighborhood slices. The key idea is to maintain a slowly updating copy of the target encoder, ensuring that neighborhood features remain consistent over training while still being distinguishable from the target features.
Formally, let $\bm{\theta}_{1}$ denote the parameters of the target slice encoder, updated via standard back-propagation. We initialize the neighborhood slice encoder with parameters $\bm{\theta}_{2}=\bm{\theta}_{1}$, and thereafter update it at each iteration as:
@@
where $m\in\left[0,1\right)$ is a momentum coefficient ($0.1$ by default). We analyze the impact of $m$ in Section 5.3, finding that a relatively small momentum yields the best results.
This momentum update strikes a balance between feature consistency and distinguishability. Specifically, neighborhood slice features are extracted by an encoder whose parameters update smoothly based on the target slice encoder, reducing abrupt shifts in feature space. At the same time, because $\bm{\theta}_{2}$ lags slightly behind $\bm{\theta}_{1}$, neighborhood features remain systematically distinct from target features, helping the fusion module to distinguish intra-slice and inter-slice information more effectively.